Can someone please solve using the substitution method?
6x+6y=30
-8x+y=41

Answer by richwmiller(17219)

divide the first by 6
x+y=5
-8x+y=41
y=5-x
-8x+5-x=41
-9x=36
x=-4
x+y=5
-4+y=5
y=9
(-4,9)